OpenStudy (anonymous):

How to do this sum? suppose that f(x)=dx+m,(d&m are real numbers). Given that f(f(f(x))) = 8x+21, what is the value of d+m? Please help me or give me a hint.

hartnn (hartnn):

f(x)=dx+m can u find f(f(x)) =... ?

hartnn (hartnn):

replace 'x' in f(x) by 'dx+m'

OpenStudy (anonymous):

f(f(x))= d^2x +dm+ m right? nd f(f(f(x)))=d^3x + d^2m + dm + m

hartnn (hartnn):

yes.

hartnn (hartnn):

d^3x + d^2m + dm + m = 8x+21 so d^3 = 8 comparing the co-efficients of x so, d=... ?

hartnn (hartnn):

equating constants d^2m + dm + m = (d^2+d+1) m = 21

OpenStudy (anonymous):

@hartnn thank u so much!!!!

hartnn (hartnn):

welcome ^_^